OpenStudy (anonymous):

1/5 of the students of a certain high school does not attend flag ceremony. If 5 students are selected, what is the probability that at least 1 did not attend the flag ceremony?

OpenStudy (shubhamsrg):

req probab = 1- P(all attended)

OpenStudy (anonymous):

The probability all attended is \(\left(\frac45\right)^5\). The probability at least 1 did not attend is then \(1-\left(\frac45\right)^5\).

OpenStudy (anonymous):

So is the asnwer 2101/3125?

OpenStudy (jack1):

.67issh so 67 percent...?